A wave traveling 87 m/s has a wavelength of 4.6 m. What is the frequency of the wave (in Hz)?



1
Expert's answer
2021-08-16T08:37:16-0400

We can find the frequency of the wave from the wave speed formula:


"v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{87\ \dfrac{m}{s}}{4.6\ m}=18.9\ Hz."